OpenStudy (bekkah323):

factor: 2x^3+6x^2+x+3

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

\[(2x^{2}+1)(x+3)\]

OpenStudy (bekkah323):

how did you get there @etnoggin

OpenStudy (anonymous):

sorry bout that.. i was wrong i didnt see the factor word LoL

OpenStudy (anonymous):

(x + 3)(2x^2 + 1)

OpenStudy (bekkah323):

how did you get there @TjRoDz

OpenStudy (anonymous):

try using foil method on either my equation or etnoggin's..

OpenStudy (bekkah323):

i did and it put it back into the original form, u

OpenStudy (anonymous):

 hmm lets see (x)(2x^2) = 2x^3 (3)(1) = 3 (3)(2x^2) = 6x^2 (1)(x) = 1x or just x lemme combine 2x^3 + 3 + 6x^2 + x so it matches :)